# From Symmetry to Structure: Gauge-Invariant Operators in Multi-Matrix Quantum Mechanics

### v1

@@ B.1 $f=1$ and $b_{1}=0=b_{2}$
Tr (($\mbox{Tr\,}\left((\Gamma_{12}\Gamma_{21})^{n}\right)$ (101)
@@
×$\times\frac{1}{(1-{t_{11}w_{12}\over t_{21}t_{22}})(1-{t_{11}t_{12}w_{12}\over t_{21}t_{22}})(1-{t_{21}w_{21}\over t_{11}})(1-{t_{21}w_{21}\over t_{11}t_{12}})}$
×$\times\frac{1}{(1-{t_{21}t_{22}w_{21}\over t_{11}})(1-{t_{21}t_{22}w_{21}\over t_{11}t_{12}})}$
@@ B.3 $N=2$ and $b_{1}=0=b_{2}$
As discussed in the main text, the total number of primary invariants is given by $18(f-1)+$ 1. For this case we can do even better and we find that, for $f>1$, there are $4(f-1)$ primaries consisting of one pair of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals and $14(f-1)+$ 1 primaries consisting of two pairs of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals. The number of secondary invariants grows exponentially with $f$ as can be seen from Table 1).
@@ B.4 $N=3$ and $b_{1}=0=b_{2}$
Turning now to $N=3$ the expressions for the partition functions become more involved. As with the $N=2$ case we are able to extract more detail of the structure of primary invariants. The total number of primary invariants is given by $118(f-1)+$ 1 of which (for $f>0$) we find that $3(f-1)$ consist of one pair of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals, $9(f-1)$ consist of two pairs of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals and $16(f-1)+$ 1 consist of three pairs of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals. The number of secondary invariants grow rapidly with $f$ as can be seen from Table 2.
@@ Appendix C Secondary Invariants at $N=2$
In this case there are $d$ primary invariants of length 1 and $33d-$ 3 primary invariants of length 2 so that the general form of the partition function is

### v2

@@ B.1 $f=1$ and $b_{1}=0=b_{2}$
$\mbox{Tr\,}\left((\Gamma_{12}\Gamma_{21})^{n}\right)$ (111)
@@
$\times\frac{1}{(1-{t_{11}w_{12}\over t_{21}t_{22}})(1-{t_{11}t_{12}w_{12}\over t_{21}t_{22}})(1-{t_{21}w_{21}\over t_{11}})(1-{t_{21}w_{21}\over t_{11}t_{12}})}$
$\times\frac{1}{(1-{t_{21}t_{22}w_{21}\over t_{11}})(1-{t_{21}t_{22}w_{21}\over t_{11}t_{12}})}$
@@ B.3 $N=2$ and $b_{1}=0=b_{2}$
As discussed in the main text, the total number of primary invariants is given by 8(f1)+18(f-1)+1. For this case we can do even better and we find that, for $f>1$, there are 4(f1)4(f-1) primaries consisting of one pair of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals and 4(f1)+14(f-1)+1 primaries consisting of two pairs of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals. The number of secondary invariants grows exponentially with $f$ as can be seen from Table 1).
@@ B.4 $N=3$ and $b_{1}=0=b_{2}$
Turning now to $N=3$ the expressions for the partition functions become more involved. As with the $N=2$ case we are able to extract more detail of the structure of primary invariants. The total number of primary invariants is given by 18(f1)+118(f-1)+1 of which (for $f>0$) we find that 3(f1)3(f-1) consist of one pair of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals, 9(f1)9(f-1) consist of two pairs of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals and 6(f1)+16(f-1)+1 consist of three pairs of $\Gamma^{A}$, $\bar{\Gamma}^{B}$ bifundamentals. The number of secondary invariants grow rapidly with $f$ as can be seen from Table 2.
@@ Appendix C Secondary Invariants at $N=2$
In this case there are $d$ primary invariants of length 1 and 3d33d-3 primary invariants of length 2 so that the general form of the partition function is